# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ 4.7 Proof of restricted strong convexity for Huber loss function
$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho E\|R\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (40)
@@
$1\leq\|A\|_{F}\leq\|A\|_{1}.$
Without loss of generality, we can assume $\|A\|_{1}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
@@
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{1}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to some $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{2})\rceil$ and $l=1,\ldots,\lceil\log_{2}\frac{1}{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $H^{2}=9\log_{2}^{2}(\mu d)$.
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}E\|R\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{1}E\|R\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}E\|R\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
@@
$P(\mathcal{E})\leq P(\cup_{k,l}\mathcal{E}_{k,l})\leq H^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
@@
4.8 Other technical lemmas
@@
Recall the definitions of $P_{A_{0}}(\cdot)$ and $P_{A_{0}}^{\perp}(\cdot)$ defined in Section 1. Consider the following estimator:
$\bar{A}=\operatorname*{arg\,min}_{A\in\Omega}\left\{L(A)+\lambda\|A\|_{1}\right\},$
where $\Omega$ is the domain of $A$ and $L(A)$ is a convex loss function. If we take $\lambda\geq 3\|\nabla L(A_{0})\|$, then $\bar{A}-A_{0}$ satisfies
$\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
As special cases, we have the following inequalities. Take $\lambda\geq 3\|\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ in (2), then
$\|P_{A_{0}}^{\perp}(\hat{A}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}-A_{0})\|_{1}.$
Take $\lambda\geq 3\frac{\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$ in (8) , then
$\|P_{A_{0}}^{\perp}(\hat{A}_{S}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{S}-A_{0})\|_{1}.$
Take $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}\|$ in (10) , then
$\|P_{A_{0}}^{\perp}(\hat{A}_{H}-A_{0})\|_{1}\leq 2\|P_{A_{0}}(\hat{A}_{H}-A_{0})\|_{1}.$

### v3

@@ 4.2 Proof of restricted strong convexity for Huber loss function
$Z(\alpha,\rho)\leq 4c_{0}\alpha\rho\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}\alpha\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\alpha^{2}\log d}{n}.$ (19)
@@
$1\leq\|A\|_{F}\leq\|A\|_{*}.$
Without loss of generality, we can assume $\|A\|_{*}\in[1,\sqrt{\mu d^{3}}]$. Similarly, we have
@@
$\tilde{S}(k,l)=\{A;\|A\|_{w(F)}=1,2^{k-1}\leq\|A\|_{\infty}\leq 2^{k},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ 2^{l-1}\leq\|A\|_{*}\leq 2^{l}\}$
Every matrix $A$ with $\|A\|_{w(F)}=1$ belongs to some $\tilde{S}(k,l)$ with $k=\lfloor\log_{2}\frac{1}{d}\rfloor,\ldots,\lceil\frac{1}{2}\log_{2}(\mu d^{2})\rceil$ and $l=1,\ldots,\lceil\log_{2}\frac{1}{2}(\mu d^{3})\rceil$. Thus the total number of the pairs $(k,l)$ is less than $\Lambda^{2}=9\log_{2}^{2}(\mu d)$.
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$ (20)
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|\leq 4c_{0}2^{k}2^{l}\mathbb{E}\|\mathcal{R}\|+\sqrt{2}c_{1}2^{k}\sqrt{\frac{\mu\log d}{n}}+4c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}.$
@@
$\left|\dfrac{1}{n}\sum_{i=1}^{n}F_{A}(X_{i},\xi_{i})-f(A)\right|$ $\geq 16c_{0}\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+16c_{2}\frac{\mu\|A\|_{\infty}^{2}\log d}{n}$
$\geq 16c_{0}2^{k-1}2^{l-1}\mathbb{E}\|\mathcal{R}\|+2\sqrt{2}c_{1}2^{k-1}\sqrt{\frac{\mu\log d}{n}}+16c_{2}2^{2(k-1)}\frac{\mu\log d}{n}$
@@
$\mathbb{P}(\mathcal{E})\leq\mathbb{P}(\cup_{k,l}\mathcal{E}_{k,l})\leq\Lambda^{2}\exp(-2\mu\log d)\leq\frac{9\log_{2}^{2}\mu d}{d^{\mu}}\cdot\frac{1}{d^{\mu}}\leq\frac{1}{d}$
@@
4.3 Proof of Theorem 3
From Appendix A in Klopp [2014], we know the following two inequalities. Given $\lambda\geq 3\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$, we have
$\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}\leq\frac{5}{3}\lambda\sqrt{r}\|\hat{A}-A_{0}\|_{F}.$ (21)
We also have
$\|\hat{A}-A_{0}\|_{*}\leq\sqrt{72r}\|\hat{A}-A_{0}\|_{F}.$ (22)
Regarding the difference between $\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{0}\rangle^{2}$ and $\|\hat{A}-A_{0}\|^{2}_{w(F)}$, we have the following lemma.
@@ 4.3 Proof of Theorem 3
Assume $\|A\|_{w(F)}^{2}=1$. Then there is a constant $C$ dependent on $c_{0},c_{1},c_{2}$ such that
$\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A\rangle^{2}-\|A\|^{2}_{w(F)}\right|\leq C\left(\|A\|_{\infty}\|A\|_{*}\mathbb{E}\|\mathcal{R}\|+\|A\|_{\infty}\sqrt{\frac{\mu\log d}{n}}+\frac{\mu\|A\|_{\infty}^{2}\log d}{n}\right)$ (23)
for all such $A$, uniformly with probability at least $1-\frac{1}{d}$. Here $\mathcal{R}=\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}$.